Home‎ > ‎Crew 138‎ > ‎

317-EngineeringReport

Engineering Report
03/17/2014
Jay Berger
Was Power Interrupted:
no
Time Of Power Outage:

Length Of Power Outage:

Diesel Fuel Level:
0.25
Propane Fuel Level:
0.55
Gasoline Fuel Level:
1
Fuel Notes And Comments:

Atv Oil Quarts:
N.A.
Car Oil Quarts:
N.A.
Oil Notes And Comments:

Trailer Potable Water Level:
30.5
Trailer Potable Water Refilled:
no
Outside Potable Water Level:
30
Outside Potable Water Refilled:
no
Loft Potable Water Gallons:
55
Loft Potable Water Refilled:
yes
Potable Water Meter Reading:
80,192.8
Greenhouse Notes And Comments:
none
Telescope Used:
no
Observatory Notes And Comments:

Network Failed:
no
Network Throttled:
no
Network Signal Strength:
189
Network Download Allowane:
53
Network Notes And Comments:

Hvac Air Conditioner Used:
no
Hvac Furnace Used:
no
Hvac Lab Wall Heater Used:
no
Hvac Upper Deck Thermostat Setting:
66
Hvac Upper Deck Temperature Reading:
70
Hvac Lower Deck Thermostat Setting:
n/a
Hvac Lower Deck Temperature Reading:
72
Hvac Notes And Comments:

Hab Car Used:
yes
Hab Car Oil Checked:
no
Hab Car Tire Status:
Good
Hab Car Odometer:
132,340.1
Hab Car Notes And Comments:

Atv One Used:
no
Atv One Oil Checked:
no
Atv One Fuel Used Gals:

Atv One Tires Status:

Atv One Hours Used Day:

Atv One Hours Used Year:

Atv One Notes And Comments:

Atv Two Used:
no
Atv Two Oil Checked:
no
Atv Two Fuel Used Gals:

Atv Two Tires Status:

Atv Two Hours Used Day:

Atv Two Hours Used Year:

Atv Two Notes And Comments:

Atv Three Used:
no
Atv Three Oil Checked:
no
Atv Three Fuel Used Gals:

Atv Three Tires Status:

Atv Three Hours Used Day:

Atv Three Hours Used Year:

Atv Three Notes And Comments:

Atv Four Used:
no
Atv Four Oil Checked:
no
Atv Four Fuel Used Gals:

Atv Four Tires Status:

Atv Four Hours Used Day:

Atv Four Hours Used Year:

Atv Four Notes And Comments:

Atv Five Used:
no
Atv Five Oil Checked:
no
Atv Five Fuel Used Gals:

Atv Five Tires Status:

Atv Five Hours Used Day:

Atv Five Hours Used Year:

Atv Five Notes And Comments:

Comments:
HEALTH EMERGENCY:  Septic tank below is full and overflowing.  There is a stream of black water running from under hab through ATV parking and to circle drive in front of hab.  Smell downstairs in lab and especially in bathroom facilities is pungent and there is concern for toxic atmosphere.  Shannon was here on Saturday and Sunday and confirmed the overflow issue and said she would have the septic tank drained.  Please acknowledge request.

Due to cleaning activities okayed by Shannon, water use and trash quantity have been high.  Plans should be made to get trailer for water refill and trash dump in the next few days. 

Cleaned and organized EVA room.  Repaired backpacks.  All 6 backpacks are now functional.